a star is moving away from earth with speed V.The change in wavelength observed on earth is?

Grade:12

1 Answers

Aarti Gupta
10 years ago
If the star is moving away from us, its waves are effectively stretched out when they reach Earth, increasing their wavelength. This shifts the star's spectral lines toward the red end of the spectrum.
In 1 second there are f waves and these will cover a distance of v + c where c is the speed of light. (Speed = distance/time and time = 1 second)

The apparent wavelength observed on earth in line with the motion is ?a and is given by

?a = (v + c)/f but f = c/? so ?a = (v + c) ? /c = (v/c + 1) ?

so

?a - ? = v? /c differences in wavelength can be written as ?? so (?a - ? = ??) so ?? = v?/c

Thanks & Regards

Aarti gupta

askiitians Faculty